An electic fan rotating at 3 rev/s-1 is switched off.it comes to rest in 18 sec,the deceleration produced is?
1
Expert's answer
2021-03-04T17:22:28-0500

We have,

Initial angular velocity, "\omega_o" = 3 rev/s

Final angular velocity, "\omega" = 0

Time t = 18 sec

Using first equation of circular motion,

"\omega = \omega_o + \alpha t"

"0 = 3 + \alpha\times 18"

"\alpha = -0.166 rev/s^2"